# MolGround: A Benchmark for Molecular Grounding

### v2

@@ Abstract
Current molecular understanding approaches predominantly focus on the descriptive aspect of human perception, providing broad, topic-level insights. However, the referential aspect—linking molecular concepts to specific structural components—remains largely unexplored. To address this gap, we propose a molecular grounding benchmark designed to evaluate a model’s referential abilities. We align molecular grounding with established conventions in NLP, cheminformatics, and molecular science, showcasing the potential of NLP techniques to advance molecular understanding within the AI for Science movement. Furthermore, we constructed the largest molecular understanding benchmark to date, comprising 79k QA pairs, and developed a multi-agent grounding prototype as proof of concept. This system outperforms existing models, including GPT-4o, and its grounding outputs have been integrated to enhance traditional tasks such as molecular captioning and ATC (Anatomical, Therapeutic, Chemical) classification.
@@
Jiaxin Wu1, Ting Zhang1, Rubing Chen1, Wengyu Zhang1, Chen Jason Zhang1, Xiaoyong Wei2,1,thanks: Corresponding Author, Li Qing1, 1The Hong Kong Polytechnic University, 2Sichuan University,
@@ 2 Related Work
The complementary relationship between descriptive and referential perceptions has been extensively explored in cognitive science, as seen in Fregean Semantics (senses and references) Cocchiarella (1974), Russell’s Theory of Descriptions (definite descriptions and proper names) Geach (1950), and Discourse Representation Theory (descriptions and referents) Kamp and Reyle (1993). However, referential perception has not been explicitly modeled or systematically evaluated in molecular understanding. From this perspective, Table 1 summarizes advanced models such as BioT5 Pei et al. (2024), ChemLLM Zhang et al. (2024a), and Mol-Instructions Fang et al. (2024), alongside commonly adopted benchmarks like ChemBench4K Zhang et al. (2024a), and MoleculeQA Lu et al. (2024). It is clear that this is an area requiring more focused and explicit attention. While recent advancements in models have made strides toward incorporating referential perception, with promising results observed in integrating referential perception-oriented visual grounding (e.g., RefFormer Wang et al. (2024), ClawMachine Ma et al. (2024), DOrA Wu et al. (2024)), significant challenges remain. This results from the heavy reliance on costly human expertise for benchmark construction and the lack of a systematic formulation of the problem. As highlighted in Table 1 , our proposed MolGround represents an initial effort to address these challenges. It scales up to 1.28 times larger than existing benchmarks and introduces fine-grained definitions to better align with the requirements of referential perception.
@@
MolGround (ours) Name Entity (CNER) 1442 0.00 100.00
Name2Struct (BNSM) 1370 100.00 0.00
Localization (RSL) 27,824 0.00 100.00
Grounding (SRG) 12,474 0.00 100.00
Analysis (SFA) 35,898 0.00 100.00
Total 79,008 1.73 98.27
@@ 4.3 Structurization and Annotation
The structuring and annotation process is performed iteratively, allowing our grounding agent to collaborate with human chemical experts. The 79k QA pairs listed in Table 1 are constructed based on the human annotation data. We split the QAs into training, validation, and testing sets using a split ratio (80%, 10%, 10%) on each task, and ensure that there is no overlapped molecule between different sets.
@@ 5 Experiments
Metric F1 Acc. Acc. Acc. $F1_{l}$ $IoU_{l}$ $Acc_{g}$ $F1_{l}$ $IoU_{l}$ $Acc_{g}$ $Cov_{s}$
LLM GPT4o 0.633 0.125 0.803 0.337 0.015 0.148 0.755 0.012 0.059 0.311 0.399
LLaMA 3.1-8B 0.504 0.092 0.574 0.111 0.006 0.175 0.672 0.000 0.018 0.083 0.126
LLaMA 3.1-70B 0.637 0.063 0.465 0.233 0.008 0.100 0.473 0.001 0.009 0.031 0.045
@@
ChemLLM-7B 0.000 0.000 0.005 0.154 0.000 0.000 0.678 0.000 0.000 0.000 0.000
Mol-Instructions 0.152 0.000 0.233 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
MLLM GPT4o-Vision 0.578 0.246 0.558 0.321 0.004 0.052 0.332 0.000 0.001 0.012 0.016
LLaVA-Next-7B 0.412 0.021 0.088 0.142 0.020 0.174 0.737 0.000 0.001 0.004 0.005
@@ 5.1 Baselines
Metric F1 Acc. Acc. Acc. $F1_{l}$ $F1_{l}$
Baselines GPT4o 0.633 0.125 0.803 0.337 0.015 0.012
LLaMA 3.1-8B 0.504 0.092 0.574 0.111 0.006 0.000
Mol-Instructions 0.152 0.000 0.233 0.001 0.000 0.000
Baselines + ICL (Few-shot) GPT4o 0.721 0.314 0.685 0.754 0.017 0.270
LLaMA 3.1-8B 0.722 0.180 0.670 0.587 0.003 0.120
Mol-Instructions 0.350 0.000 0.276 0.117 0.000 0.036
Baselines + ICL (RAG) GPT4o 0.915 0.397 0.444 0.754 0.174 0.171
LLaMA 3.1-8B 0.881 0.361 0.162 0.655 0.149 0.113
Mol-Instructions 0.864 0.083 0.059 0.561 0.091 0.072
Baselines + SFT LLaMA 3.1-8B 0.641 0.426 0.602 0.899 0.275 0.315
Mol-Instructions 0.727 0.397 0.604 0.917 0.295 0.337
@@ 5.2 Evaluation of Pretrained Models
Table 3 compares the performance of LLMs and MLLMs across five molecular grounding tasks. Overall, most tasks remain challenging for all baseline models, with accuracy generally below 0.5. In particular, BNSM, SFA, and RSL prove to be the most difficult, with all models achieving accuracies below 0.337. By contrast, CNER and SRG exhibit relatively better performance. The highest F1-score for CNER is 0.633 and the highest accuracy for SRG is 0.803, which are achieved by GPT-4o. For the CNER task, models often over-extract functionality words as chemical substructure names (e.g., “amide functionality”) and struggle to correctly assign roles to extracted chemical names. For the SFA, models perform well on counting non-existent rings but struggle with monocyclic ring identification due to subtle structural differences. For example, the predicted probabilities of the benzene to be a donor or acceptor is Moreover, BioT5+ fails to follow instructions, and Mol-Instructions exhibits a strong bias in selecting question choices . While chemical name recognition (CNER) is relatively well-handled, with most models achieving an F1-score above 0.9, smaller models (e.g., LLaMA 3.1-8B, ChemLLM-7B) similar names (e.g., ether vs. ester) For more complex tasks (BNSM and SRG), performance remains low, even for the best model, GPT-4o, which achieves only 0.125 and 0.333 accuracy, respectively. In BNSM, models perform well for simple structures but struggle with complex mappings, often generating incorrect yet structurally similar names or SMILES strings. SRG results indicate difficulty in establishing substructures’ relationships, with models misinterpreting textual cues and neglecting structural connections. RSL task witnessed a particularly poor performance, with the best F1 (0.020) and IoU (0.174). In singular RSL, GPT-4o achieves the highest accuracy (0.755) but with low IoU (0.148), while LLaMA 3.1-8B prioritizes broader highlighting at the cost of accuracy. For multiple RSL, most models generate invalid outputs and show low substructure coverage (max 39.9%). MLLMs, despite access to structural images, do not surpass LLMs (except BNSM).
@@ 6 Conclusion
This paper has introduced a molecular grounding benchmark to enhance the referential aspect of molecular understanding. MolGround, with 79k QA pairs across five subtasks, is the largest molecular QA benchmark. Our evaluation shows that existing LLMs struggle with these tasks, with SFT and ICL yielding minor improvements. Our grounding prototype outperforms existing models and enhances downstream tasks like captioning and ATC classification.

### v4

@@ Abstract
Current molecular understanding approaches predominantly focus on the descriptive aspect of human perception, providing broad, topic-level insights. However, the referential aspect—linking molecular concepts to specific structural components—remains largely unexplored. To address this gap, we propose a molecular grounding benchmark designed to evaluate a model’s referential abilities. We align molecular grounding with established conventions in NLP, cheminformatics, and molecular science, showcasing the potential of NLP techniques to advance molecular understanding within the AI for Science movement. Furthermore, we constructed the largest molecular understanding benchmark to date, comprising 117k QA pairs, and developed a multi-agent grounding prototype as proof of concept. This system outperforms existing models, including GPT-4o, and its grounding outputs have been integrated to enhance traditional tasks such as molecular captioning and ATC (Anatomical, Therapeutic, Chemical) classification.
@@
Jiaxin Wu1, Ting Zhang1, Rubing Chen1, Wengyu Zhang1, Chen Jason Zhang1, Xiao-Yong Wei2,1,thanks: Corresponding Author, Li Qing1, 1The Hong Kong Polytechnic University, 2Sichuan University,
@@ 2 Related Work
The complementary relationship between descriptive and referential perceptions has been extensively explored in cognitive science, as seen in Fregean Semantics (senses and references) Cocchiarella (1974), Russell’s Theory of Descriptions (definite descriptions and proper names) Geach (1950), and Discourse Representation Theory (descriptions and referents) Kamp and Reyle (1993). However, referential perception has not been explicitly modeled or systematically evaluated in molecular understanding. From this perspective, Table 1 summarizes advanced models such as BioT5 Pei et al. (2024), ChemLLM Zhang et al. (2024a), and Mol-Instructions Fang et al. (2024), alongside commonly adopted benchmarks like ChemBench4K Zhang et al. (2024a), and MoleculeQA Lu et al. (2024). It is clear that this is an area requiring more focused and explicit attention. While recent advancements in models have made strides toward incorporating referential perception, with promising results observed in integrating referential perception-oriented visual grounding (e.g., RefFormer Wang et al. (2024), ClawMachine Ma et al. (2024), DOrA Wu et al. (2024)), significant challenges remain. This results from the heavy reliance on costly human expertise for benchmark construction and the lack of a systematic formulation of the problem. As highlighted in Table 1 , our proposed MolGround represents an initial effort to address these challenges. It scales up to 1.91 times larger than existing benchmarks and introduces fine-grained definitions to better align with the requirements of referential perception.
@@
MolGround (ours) Name Entity (CNER) 15,033 0.00 100.00
Name2Struct (BNSM) 1,397 100.00 0.00
Localization (RSL) 48,107 0.00 100.00
Grounding (SRG) 13,994 0.00 100.00
Analysis (SFA) 38,780 0.00 100.00
Total 117,311 1.19 98.81
@@ 4.3 Structurization and Annotation
The structuring and annotation process is performed iteratively, allowing our grounding agent to collaborate with human chemical experts. The 117k QA pairs listed in Table 1 are constructed based on the human annotation data. We split the QAs into training, validation, and testing sets using a split ratio (80%, 10%, 10%) on each task, and ensure that there is no overlapped molecule between different sets.
@@ 5 Experiments
Metric F1 Acc. F1 Acc. $F1_{l}$ $IoU_{l}$ $Acc_{g}$ $F1_{l}$ $IoU_{l}$ $Acc_{g}$ $Cov_{s}$
LLM GPT-4o 0.633 0.125 0.160 0.337 0.015 0.148 0.755 0.011 0.073 0.429 0.543
LLaMA 3.1-8B 0.504 0.092 0.001 0.111 0.006 0.175 0.672 0.000 0.018 0.083 0.126
LLaMA 3.1-70B 0.637 0.063 0.004 0.233 0.008 0.100 0.473 0.001 0.009 0.031 0.045
@@
ChemLLM-7B 0.000 0.000 0.000 0.154 0.000 0.000 0.678 0.000 0.000 0.000 0.000
Mol-Instructions 0.152 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
MLLM GPT-4o-Vision 0.578 0.246 0.061 0.321 0.004 0.052 0.332 0.000 0.001 0.012 0.016
LLaVA-Next-7B 0.412 0.021 0.042 0.142 0.020 0.174 0.737 0.000 0.001 0.004 0.005
@@ 5.1 Baselines
Metric F1 Acc. F1 Acc. $F1_{l}$ $F1_{l}$
Baselines GPT-4o 0.633 0.125 0.160 0.337 0.015 0.011
LLaMA 3.1-8B 0.504 0.092 0.001 0.111 0.006 0.000
Mol-Instructions 0.152 0.000 0.000 0.001 0.000 0.000
Baselines + ICL (Few-shot) GPT-4o 0.721 0.314 0.385 0.754 0.017 0.270
LLaMA 3.1-8B 0.722 0.180 0.195 0.587 0.003 0.120
Mol-Instructions 0.350 0.000 0.000 0.117 0.000 0.036
Baselines + ICL (RAG) GPT-4o 0.915 0.397 0.368 0.754 0.174 0.171
LLaMA 3.1-8B 0.881 0.361 0.072 0.655 0.149 0.113
Mol-Instructions 0.864 0.083 0.000 0.561 0.091 0.072
Baselines + SFT LLaMA 3.1-8B 0.641 0.426 0.078 0.899 0.275 0.315
Mol-Instructions 0.727 0.397 0.072 0.917 0.295 0.337
@@ 5.2 Evaluation of Pretrained Models
Table 3 compares the performance of LLMs and MLLMs across five molecular grounding tasks. Overall, most tasks remain challenging for all baseline models, with accuracy generally below 0.5. In particular, BNSM, SFA, and RSL prove to be the most difficult, with all models achieving accuracies below 0.337. By contrast, CNER and SRG exhibit relatively better performance. The highest F1-score for CNER is 0.633 and the highest accuracy for SRG is 0.803, which are achieved by GPT-4o. For the CNER task, models often over-extract functionality words as chemical substructure names (e.g., “amide functionality”) and struggle to correctly assign roles to extracted chemical names. For the SFA, models perform well on counting non-existent rings but struggle with monocyclic ring identification due to subtle structural differences. For example, the predicted probabilities of the benzene to be a donor or acceptor is Moreover, BioT5+ fails to follow instructions, and Mol-Instructions exhibits a strong bias in selecting question choices . While chemical name recognition (CNER) is relatively well-handled, with most models achieving an F1-score above 0.9, smaller models (e.g., LLaMA 3.1-8B, ChemLLM-7B) similar names (e.g., ether vs. ester) For more complex tasks (BNSM and SRG), performance remains low, even for the best model, GPT-4o, which achieves only 0.125 and 0.333 accuracy, respectively. In BNSM, models perform well for simple structures but struggle with complex mappings, often generating incorrect yet structurally similar names or SMILES strings. SRG results indicate difficulty in establishing substructures’ relationships, with models misinterpreting textual cues and neglecting structural connections. RSL task witnessed a particularly poor performance, with the best F1 (0.020) and IoU (0.174). In singular RSL, GPT-4o achieves the highest accuracy (0.755) but with low IoU (0.148), while LLaMA 3.1-8B prioritizes broader highlighting at the cost of accuracy. For multiple RSL, most models generate invalid outputs and show low substructure coverage (max 54.3%). MLLMs, despite access to structural images, do not surpass LLMs (except BNSM).
@@ 6 Conclusion
This paper has introduced a molecular grounding benchmark to enhance the referential aspect of molecular understanding. MolGround, with 117k QA pairs across five subtasks, is the largest molecular QA benchmark. Our evaluation shows that existing LLMs struggle with these tasks, with SFT and ICL yielding minor improvements. Our grounding prototype outperforms existing models and enhances downstream tasks like captioning and ATC classification.